# What Makes a Maze Look Like a Maze?

### v1

@@ Abstract
A unique aspect of human visual understanding is the ability to flexibly interpret abstract concepts: acquiring lifted rules explaining what they symbolize, grounding them across familiar and unfamiliar contexts, and making predictions or reasoning about them. While off-the-shelf vision-language models excel at making literal interpretations of images (e.g., recognizing object categories such as tree branches), they still struggle to make sense of such visual abstractions (e.g., how an arrangement of tree branches may form the walls of a maze). To address this challenge, we introduce Deep Schema Grounding (DSG), a framework that leverages explicit structured representations of visual abstractions for grounding and reasoning. At the core of DSG are schemas—dependency graph descriptions of abstract concepts that decompose them into more primitive-level symbols. DSG uses large language models to extract schemas, then hierarchically grounds concrete to abstract components of the schema onto images with vision-language models. The grounded schema is used to augment visual abstraction understanding. We systematically evaluate DSG and different methods in reasoning on our new Visual Abstractions Dataset, which consists of diverse, real-world images of abstract concepts and corresponding question-answer pairs labeled by humans. We show that DSG significantly improves the abstract visual reasoning performance of vision-language models, and is a step toward human-aligned understanding of visual abstractions.
@@ 1 Introduction
Humans possess the remarkable ability to flexibly acquire and apply abstract concepts when interpreting the concrete world around us. Consider the concept “maze”: our mental model can interpret mazes constructed with conventional materials (e.g., drawn lines) or unconventional ones (e.g., icing), and reason about mazes across a wide range of configurations and environments (e.g., in a cardboard box or on a knitted square). Our goal is to build systems that can make such flexible and broad generalizations as humans do. This necessitates a reconsideration of a fundamental question: what makes a maze look like a maze? A maze is not defined by concrete visual features such as the specific material of walls or its perpendicular intersections, but by lifted rules over symbols—a plausible model for a maze includes its layout, the materials composing the walls, and the designated entry and exit. Crucially, these model components can be realized by infinitely many real-world variations. For instance, walls constructed from candy canes, hay, or popsicle sticks, despite their diverse materials, are all identifiable as parts of a maze.
@@
However, current vision-language models (VLMs) often struggle to reason about visual abstractions at a human level, frequently defaulting to literal interpretations of images, such as a collection of object categories. These interpretations may be technically correct descriptions of the scene, but may not align with the abstract concept underlying the image. For example, when given the intended concept of “maze” and an image that realizes the concept with unconventional objects, VLMs such as GPT-4V (OpenAI, 2023) and LLaVA (Liu et al., 2024) fail to correctly recognize objects in the scene as components of the visual abstraction. It is unclear whether these off-the-shelf models leverage knowledge of abstract concepts to make sense of images as humans do.
@@
To address the challenge of understanding abstract visual concepts, we propose Deep Schema Grounding (DSG), a framework for models to interpret visual abstractions. At the core of DSG are schemas—a dependency graph description of abstract concepts (see Figure 1). Schemas characterize common patterns that humans use to interpret and reason about the visual world (Schank & Abelson, 1975). A schema for “helping” allows us to understand relations between characters in a finger puppet scene, while a schema for “tic-tac-toe” allows us to play the game even when the grid is composed of hula hoops instead of drawn lines. A schema for “maze” is what makes a maze look like a maze.
@@
DSG explicitly uses schemas generated by and grounded by large pretrained models to reason about visual abstractions. Concretely, we model schemas as programs encoding directed acyclic graphs (DAGs), which decompose an abstract concept into a set of more concrete visual concepts as subcomponents, as illustrated in Figure 1. The full framework is composed of three steps. First, we extract schema definitions of abstract concepts from a large language model (LLM). Next, DSG hierarchically queries a vision-language model (VLM), first grounding concrete symbols in the DAG (i.e. symbols that do not depend on the interpretation of other symbols), then using those symbols as conditions to ground more abstract symbols. Finally, we use the resolved schema, including the grounding of all its components, as an additional input into a vision-language model to provide holistic context about the image, such that we can improve visual reasoning across diverse downstream queries about the abstract concept. Our method is a general framework for abstract concepts that does not depend on specific models; the LLMs and VLMs used are interchangeable.
@@
To investigate the capabilities of models in understanding visual abstractions, we introduce the Visual Abstractions Dataset (VAD). VAD is a visual question-answering dataset that consists of diverse, real-world images representing abstract concepts. The abstract concepts span $4$ different categories: strategic concepts that are characterized by rules and patterns (e.g., “tic-tac-toe”), scientific concepts of phenomena that cannot be visualized in their canonical forms (e.g., “cell”), social concepts that are defined by theory-of-mind relations (e.g., “deceiving”), and domestic concepts of household objectives that cannot be directly defined by specific arrangements of objects (e.g, “table setting for two”). Each image is an instantiation of an abstract concept, and is paired with questions that probe understanding of the visual abstraction; for example, “Imagine that the image represents a maze. Are there clear entrances or exits in the maze?” The Visual Abstractions Dataset comprises $540$ of such examples, with answers labeled by $5$ human annotators from Prolific. We found that off-the-shelf VLMs and integrated LLMs with APIs have significant room for improvement on this benchmark.
@@
We evaluate Deep Schema Grounding on the Visual Abstractions Dataset, and show that DSG consistently improves performance of vision-language models across question types, abstract concept categories, and base models. Notably, DSG improves GPT-4V by $5.4$ percent point overall ($\uparrow 8.3\%$ relative improvement), and, in particular, demonstrates a $6.7$ percent point improvement ($\uparrow 11.0\%$ relative improvement) in questions that involve counting. While the challenge of visual abstraction understanding is still far from being fully solved, our results show that DSG is a promising solution that effectively schemas to structure the thinking process of visual reasoning systems.
@@
  • We introduce the Visual Abstractions Dataset (VAD), a diverse, real-world visual question-answering benchmark that evaluates VLMs’ understanding of visual abstractions.
@@ 2 Related Works
Our approach is related to recent works on chain-of-thought reasoning in vision-language models (Lu et al., 2022; Chen et al., 2023), hierarchical image classification (Koh et al., 2020; Menon & Vondrick, 2022), hierarchical relation detection (Li et al., 2024), and visual reasoning via programs (Hu et al., 2017; Mao et al., 2019; Hsu et al., 2024). In contrast to these algorithms, which are primarily designed for object and relational reasoning, we focus on the grounding of scene-level concepts. Our framework also shares a similar spirit to prior approaches (Menon & Vondrick, 2022; Li et al., 2024) that leverage vision-language models to generate text descriptions of class features for classification tasks. However, our paper proposes the generation of hierarchical schemas with LLMs to ground visual abstractions for reasoning. Additionally, our hierarchical schema representation is related to hierarchical program representations used in text-only reasoning (Gao et al., 2023; Zelikman et al., 2023; Zhang et al., 2023); in contrast, we leverage program representations for understanding concepts in natural images. Notably, DSG performs schema inference with a VLM; compared to fully symbolic inference processes, our framework is capable of leveraging all visual information in the image for interpreting abstractions.
@@ 3.1 Visual Abstraction Schema
The most important feature of a visual abstraction schema is that the definition of such concepts is universal, in the sense that the program describes a universal characterization of mazes made of various components, rather than being restricted to specific visual instantiations (e.g., mazes formed by drawn lines). Schemas enable machines to better generalize to novel scenes.
@@ 3.2 Extracting Schemas of Abstract Concepts
In particular, our goal is to extract a universal schema for abstract concepts that is not instance-specific, but is applicable across diverse visual stimuli. Notably, our prompt to the LLMs is generic and simple, with only one example of a schema for the abstract concept “academia” (which consists of key components: faculty, students, etc.), and a short instruction of the schema generation task. We find that LLMs demonstrate a remarkable ability to extrapolate across a wide array of concepts, and produce faithful decompositions of the visual abstractions. We detail the prompt and LLM-generated schemas for each concept in the Appendix.
@@ 3.3 Hierarchically Grounding Schemas on Images
Grounding components following the hierarchical schema structure is crucial for abstract concepts with many possible instantiations. For example, for the “maze” concept, DSG first resolves components that are more coarse and concrete (e.g., layout), then uses those components as conditions to resolve more difficult grounding problems of fine-grained and abstract components (e.g., entry-exit). Figure 3 shows examples of schemas and how they may be grounded onto images. More examples and failure modes can be found in the Appendix.
@@ 3.4 Visual Question-Answering Augmented with Grounded Schemas
4 Visual Abstractions Dataset
In order to evaluate models on visual abstraction reasoning, we propose a new dataset, the Visual Abstractions Dataset (VAD), illustrated with examples in Figure 5. It consists of $180$ images and $3$ questions per image, with a total of $540$ test examples. VAD is comprised of $12$ different abstract concepts spanning $4$ categories, where each concept is associated with $15$ examples of real-world images. The questions bridge multiple question types: binary-choice, counting, and open-ended questions. We detail the set of abstract concepts and the image and text components of VAD below.
Figure 5: The Visual Abstractions Dataset comprises diverse, real-world images that represent $12$ different abstract concepts across $4$ categories.
@@ 4 Visual Abstractions Dataset
We create $3$ types of natural language questions for each abstract concept in VAD. These questions ask about different aspects of the images and take various forms, including binary-choice questions, counting questions, and general open-ended questions. In total, there are $6$ types of counting questions, $14$ types of binary-choice questions, and $16$ types of open-ended questions. An example of a binary-choice question is “Imagine that the image represents table setting for two. What is the configuration of the seats? Describe the answer as across or adjacent.”; an example of a counting question is “Imagine that the image represents a cell. Excluding the nucleus, how many types of organelles are there?”; an example of an open-ended question is “Imagine that the image represents a treasure map. Based on the image, find the starting location; what is the main color of the map region that is physically closest to the starting location?” The overall dataset comprises $540$ of such questions.
Answers to all image-question pairs are labeled by $5$ human annotators from Prolific; we take their modal response as the answer, to be used for exact-match text accuracy. While the Visual Abstractions Dataset focuses on multiple-choice questions and answers, we also provide a free-response variant of VAD. We present examples for each type of question along with corresponding images and answers in the Appendix, and release our dataset here.
@@ 5 Results
Table 1: DSG outperforms prior works across all questions and different question types in the Visual Abstractions Dataset. Results are averaged over $5$ runs.
@@
GPT-4V (OpenAI, 2023) $0.653$ $0.611$ $0.659$ $0.663$
DSG (ours) $\mathbf{0.707}$ $\mathbf{0.678}$ $\mathbf{0.707}$ $\mathbf{0.719}$
@@
Table 2: DSG improves the performance of GPT-4V across categories of abstract concepts.
@@
GPT-4V (OpenAI, 2023) $0.644$$\pm 0.033$ $0.572$$\pm 0.098$ $0.679$$\pm 0.098$ $0.717$$\pm 0.053$
DSG (ours) $\mathbf{0.721}$$\pm 0.057$ $\mathbf{0.631}$$\pm 0.134$ $\mathbf{0.724}$$\pm 0.071$ $\mathbf{0.753}$$\pm 0.058$
We validate Deep Schema Grounding in comparison to baseline VLMs and integrated LLMs with APIs on the Visual Abstractions Dataset. We first compare our work with ViperGPT (Surís et al., 2023), VisProg (Gupta & Kembhavi, 2023), LLaVA (Liu et al., 2024), InstructBLIP (Dai et al., 2024), and GPT-4V (OpenAI, 2023). Then, we ablate different components of the DSG framework. Finally, we show that DSG is model-agnostic and can benefit both closed-source and open-source models. Baseline models are given the abstract concept itself in the prompt along with the question (e.g., “Imagine that the image represents a maze.”), but without DSG’s resolved schema of the concept. All results are averaged over $5$ runs. Here, we focus on a systematic evaluation of multiple-choice answers, where accuracy is calculated as exact-match accuracy. In the Appendix, we provide additional comparisons with a graded accuracy metric based on the degree of alignment to human judgments, and report results on free-form answers evaluated by BERTScores (Zhang et al., 2020).
@@
Table 1 summarizes the multiple-choice accuracy of DSG with GPT-4V as the base VLM in comparison to that of prior works across $5$ runs. DSG significantly outperforms VLMs and integrated LLMs with APIs across all questions and all types of questions. In particular, our method demonstrates a boost in accuracy of $5.4$ percent point overall ($\uparrow 8.3\%$ relative improvement) compared to GPT-4V. Across different question types, DSG shows the largest performance improvement of $6.7$ percent point ($\uparrow 11.0\%$ relative improvement) for counting-based questions, such as “How many entities are helping?” We hypothesize that the significant improvement in accuracy for these questions is due to more accurate labels for objects with DSG, as the schema grounding step will concretely associate concepts such as “the entity who is helping” with particular parties in the image (e.g., pandas). We see that these questions usually require a more sophisticated reasoning process to answer correctly. In comparison, for binary-choice questions, there may exist more superficial correlations that off-the-shelf VLMs can rely on. Importantly, we also report a random human annotator’s accuracy of adhering to the modal human response, to highlight the difficulty of problems in VAD.
In Table 2, we more closely analyze the performance improvement of DSG on GPT-4V across concept categories. We report the average performance and the standard deviation for $5$ runs. We see that DSG outperforms GPT-4V across the four categories. In particular, our method demonstrates a strong boost in accuracy of $7.7$ percent point ($\uparrow 12.0\%$ relative improvement) for strategic concepts, and $5.9$ percent point ($\uparrow 10.3\%$ relative improvement) for scientific concepts. In the Appendix, we include qualitative examples of cases where DSG succeeds and fails in correctly grounding schemas. Overall, we see that DSG’s decomposition-based framework improves performance of VLMs.
@@
+ schema $0.647$ $0.613$ $0.647$ $0.659$
+ grounding $0.671$ $0.653$ $0.674$ $0.676$
+ hierarchy $0.683$ $0.651$ $0.690$ $0.688$
DSG (ours) $\mathbf{0.707}$ $\mathbf{0.678}$ $\mathbf{0.707}$ $\mathbf{0.719}$
@@
We present ablation studies of different modules in DSG in Table 3. In particular, we evaluate (1) a variant of DSG that is given only the LLM-generated schema (+ schema), without explicit grounding of each symbol, (2) a variant where DSG performs decomposed grounding sequentially instead of hierarchically (+ grounding), and (3) a hierarchical variant (+ hierarchical) where in the final question-answering step, we provide the model with only the grounding result of each schema component, instead of a full history of its grounding process.
With the first variant, we test whether GPT-4V can use the schema appropriately without a structured thinking process. We explore the second variant to test whether decomposing the visual abstraction into a set of key elements is sufficient for improving the understanding of abstract concepts, and evaluate the role of hierarchical dependencies in lower-level concepts for grounding. Finally, we present results on the third variant to test whether the full grounding trace of DSG is helpful during the final question-answering step. The results show that all these design choices help: a schema only does not lend itself to improved performance, and hierarchical grounding from concrete to abstract symbols, as well as the inclusion of the full context, improves performance in many concepts.
DSG improves other (open-source) VLMs too.
Tables 4 and 5 summarize the results of DSG with open-source LLaVA as the base VLM. DSG yields consistent improvement overall ($\uparrow 7.0\%$ relative improvement). However, in contrast to GPT-4V, both LLaVA and DSG yield poor results on counting questions, as the base VLM struggles at enumeration regardless of its abstraction understanding.

### v2

@@ Abstract
A unique aspect of human visual understanding is the ability to flexibly interpret abstract concepts: acquiring lifted rules explaining what they symbolize, grounding them across familiar and unfamiliar contexts, and making predictions or reasoning about them. While off-the-shelf vision-language models excel at making literal interpretations of images (e.g., recognizing object categories such as tree branches), they still struggle to make sense of such visual abstractions (e.g., how an arrangement of tree branches may form the walls of a maze). To address this challenge, we introduce Deep Schema Grounding (DSG), a framework that leverages explicit structured representations of visual abstractions for grounding and reasoning. At the core of DSG are schemas—dependency graph descriptions of abstract concepts that decompose them into more primitive-level symbols. DSG uses large language models to extract schemas, then hierarchically grounds concrete to abstract components of the schema onto images with vision-language models. The grounded schema is used to augment visual abstraction understanding. We systematically evaluate DSG and different methods in reasoning on our new Visual Abstractions Benchmark, which consists of diverse, real-world images of abstract concepts and corresponding question-answer pairs labeled by humans. We show that DSG significantly improves the abstract visual reasoning performance of vision-language models, and is a step toward human-aligned understanding of visual abstractions.
@@ 1 Introduction
Humans possess the remarkable ability to flexibly acquire and apply abstract concepts when interpreting the concrete world around us. Consider the concept “maze”: our mental model can interpret mazes constructed with conventional materials (e.g., drawn lines) or unconventional ones (e.g., icing), and reason about mazes across a wide range of configurations and environments (e.g., in a cardboard box or on a knitted square). Our goal is to build systems that can make such flexible and broad generalizations as humans do. This necessitates a reconsideration of a fundamental question: what makes a maze look like a maze? A maze is not defined by concrete visual features such as the specific material of walls or particular perpendicular intersections, but by lifted rules over symbols—a plausible model for a maze includes its layout, the walls, and the designated entry and exit. Crucially, lifted model components can be realized by infinitely many real-world variations, as opposed to grounded concepts which are tied to specific visual inputs. For instance, walls constructed from candy canes, hay, or popsicle sticks, despite their diverse materials, are all identifiable as parts of a maze.
@@
However, current vision-language models (VLMs) often struggle to reason about visual abstractions at a human level, frequently defaulting to literal interpretations of images, such as a collection of object categories. These interpretations may be technically correct descriptions of the scene, but may not align with the abstract concept underlying the image. For example, when given the intended concept of “maze” and an image that realizes the concept with unconventional objects, VLMs such as GPT-4o (OpenAI, 2024) and LLaVA (Liu et al., 2024) fail to correctly recognize objects in the scene as components of the visual abstraction. It is unclear whether these off-the-shelf models leverage knowledge of abstract concepts to make sense of images as humans do.
@@
To address the challenge of understanding abstract visual concepts, we propose Deep Schema Grounding (DSG), a framework for models to interpret visual abstractions. At the core of DSG are schemas—a dependency graph description of abstract concepts (see Figure 1). Schemas characterize common patterns that humans use to interpret the visual world, generalize efficiently from limited data, and reason across multiple levels of abstraction for flexible adaptation (Schank & Abelson, 1975). A schema for “helping” allows us to understand relations between characters in a finger puppet scene, while a schema for “tic-tac-toe” allows us to play the game even when the grid is composed of hula hoops instead of drawn lines. A schema for “maze” makes a maze look like a maze.
@@
DSG explicitly uses schemas generated by and grounded by large pretrained models to reason about visual abstractions. Concretely, we model schemas as programs encoding directed acyclic graphs (DAGs), which decompose an abstract concept into a set of more concrete visual concepts as subcomponents, as illustrated in Figure 1. The full framework is composed of three steps. First, we extract schema definitions of abstract concepts from a large language model (LLM). Next, DSG hierarchically queries a vision-language model (VLM), first grounding concrete symbols in the DAG (i.e. symbols that do not depend on the interpretation of other symbols), then using those symbols as conditions to ground more abstract symbols. Finally, we use the resolved schema, including the grounding of all its components, as an additional input into a vision-language model to provide holistic context about the image, such that we can improve visual reasoning across diverse downstream queries about the abstract concept. Our method is a general framework for abstract concepts without dependency on specific models; the LLMs and VLMs used are interchangeable.
@@
To investigate the capabilities of models in understanding visual abstractions, we introduce the Visual Abstractions Benchmark (VAB). VAB is a visual question-answering benchmark that consists of diverse, real-world images representing abstract concepts. The abstract concepts span $4$ different categories: strategic concepts that are characterized by rules and patterns (e.g., “tic-tac-toe”), scientific concepts of phenomena that cannot be visualized in their canonical forms (e.g., “cell”), social concepts that are defined by theory-of-mind relations (e.g., “deceiving”), and domestic concepts of household objectives that cannot be directly defined by specific arrangements of objects (e.g, “table setting for two”). Each image is an instantiation of an abstract concept, and is paired with questions that probe understanding of the visual abstraction; for example, “Imagine that the image represents a maze. Are there clear entrances or exits in the maze?” The Visual Abstractions Benchmark comprises $540$ of such examples, with answers labeled by $5$ human annotators from Prolific. We found that off-the-shelf VLMs and integrated LLMs with APIs have much room for improvement on this benchmark.
@@
We evaluate Deep Schema Grounding on the Visual Abstractions Benchmark, and show that DSG consistently improves performance of vision-language models across question types, abstract concept categories, and base models. Notably, DSG improves GPT-4o by $6.6$ percent points overall ($\uparrow 9.9\%$ relative improvement), and, in particular, demonstrates a $10$ percent point improvement ($\uparrow 16.6\%$ relative improvement) in questions that involve counting. While the challenge of visual abstraction understanding is still far from being fully solved, our results show that DSG is a promising solution that effectively uses schemas to structure the thinking process of visual reasoning systems.
@@
  • We introduce the Visual Abstractions Benchmark (VAB), a diverse, real-world visual question-answering benchmark that evaluates VLMs’ understanding of visual abstractions.
@@ 2 Related Works
Our approach is related to recent works on chain-of-thought reasoning in vision-language models (Lu et al., 2022; Chen et al., 2024), hierarchical image classification (Koh et al., 2020; Menon & Vondrick, 2023), hierarchical relation detection (Li et al., 2024), and visual reasoning via programs (Hu et al., 2017; Mao et al., 2019; Hsu et al., 2024). In contrast to these algorithms, which are primarily designed for object and relational reasoning, we focus on the grounding of scene-level concepts. Our framework also shares a similar spirit to prior approaches (Menon & Vondrick, 2023; Li et al., 2024) that leverage vision-language models to generate text descriptions of class features for classification tasks. However, our paper proposes the generation of hierarchical schemas with LLMs to ground visual abstractions for reasoning. Additionally, our hierarchical schema representation is related to hierarchical program representations used in text-only reasoning (Gao et al., 2023; Zelikman et al., 2023; Zhang et al., 2024); in contrast, we leverage program representations for understanding concepts in natural images. Notably, DSG performs schema inference with a VLM; compared to fully symbolic inference processes, our framework is capable of leveraging all visual information in the image for interpreting abstractions.
@@ 3.1 Visual Abstraction Schema
The most important feature of a visual abstraction schema is that the definition of such concepts is universal, in the sense that the program describes a universal characterization of mazes made of various components, rather than being restricted to specific visual instantiations (e.g., mazes formed by drawn lines). Hence, schemas enable machines to better generalize to novel scenes.
@@ 3.2 Extracting Schemas of Abstract Concepts
In particular, our goal is to extract a universal schema for abstract concepts that is not instance-specific, but is applicable across diverse visual stimuli. Notably, our prompt to the LLMs is generic and simple, with only one example of a schema for the abstract concept “academia” (which consists of key components: faculty, students, etc.), and a short instruction of the schema generation task. We find that LLMs demonstrate a remarkable ability to extrapolate across a wide array of concepts, and produce faithful decompositions of the visual abstractions. We detail the prompt and LLM-generated schemas for each concept in Appendix C.
@@ 3.3 Hierarchically Grounding Schemas on Images
Grounding components following the hierarchical schema structure is crucial for abstract concepts with many possible instantiations. For example, for the “maze” concept, DSG first resolves components that are more coarse and concrete (e.g., layout), then uses those components as conditions to resolve more difficult grounding problems of fine-grained and abstract components (e.g., entry-exit). Figure 3 shows examples of schemas and how they may be grounded onto images. More examples and failure modes can be found in Appendix A.
@@ 3.4 Visual Question-Answering Augmented with Grounded Schemas
4 Visual Abstractions Benchmark
In order to evaluate models on visual abstraction reasoning, we propose a new benchmark, the Visual Abstractions Benchmark (VAB), illustrated with examples in Figure 5. It consists of $180$ images and $3$ questions per image, with a total of $540$ test examples. VAB is comprised of $12$ different abstract concepts spanning $4$ categories, where each concept is associated with $15$ examples of real-world images. The questions bridge multiple question types: binary-choice, counting, and open-ended questions. We detail the set of abstract concepts and the image and text components of VAB below.
Figure 5: The Visual Abstractions Benchmark comprises diverse, real-world images that represent $12$ different abstract concepts across $4$ categories.
@@ 4 Visual Abstractions Benchmark
We create $3$ types of natural language questions for each abstract concept in VAB. These questions ask about different aspects of the images and take various forms, including binary-choice questions, counting questions, and general open-ended questions. In total, there are $6$ types of counting questions, $14$ types of binary-choice questions, and $16$ types of open-ended questions. An example of a binary-choice question is “Imagine that the image represents a table setting for two. What is the configuration of the seats? Describe the answer as across or adjacent.”; an example of a counting question is “Imagine that the image represents a cell. Excluding the nucleus, how many types of organelles are there?”; an example of an open-ended question is “Imagine that the image represents a treasure map. Based on the image, find the starting location; what is the main color of the map region that is physically closest to the starting location?” The full Visual Abstractions Benchmark comprises $540$ of such questions.
Answers to all image-question pairs are labeled by $5$ human annotators from Prolific; we take their modal response as the answer, to be used for exact-match text accuracy. While the Visual Abstractions Benchmark focuses on multiple-choice questions and answers, we also provide a free-response variant of VAB. We present examples for each type of question along with corresponding images and answers in Appendix B, and release our benchmark here.
@@ 5 Results
Table 1: DSG outperforms prior works across all questions and different question types in the Visual Abstractions Benchmark. Results are averaged over $5$ runs.
@@
GPT-4o (OpenAI, 2024) $0.664$ $0.604$ $0.655$ $0.693$
DSG (ours) $\mathbf{0.730}$ $\mathbf{0.704}$ $\mathbf{0.690}$ $\mathbf{0.776}$
@@
Table 2: DSG improves the performance of GPT-4o across categories of abstract concepts.
@@
GPT-4o (OpenAI, 2024) $0.653$$\pm 0.052$ $0.646$$\pm 0.074$ $0.647$$\pm 0.163$ $0.708$$\pm 0.082$
DSG (ours) $\mathbf{0.751}$$\pm 0.088$ $\mathbf{0.693}$$\pm 0.126$ $\mathbf{0.723}$$\pm 0.027$ $\mathbf{0.754}$$\pm 0.104$
We validate Deep Schema Grounding in comparison to baseline VLMs and integrated LLMs with APIs on the Visual Abstractions Benchmark. We first compare our work with ViperGPT (Surís et al., 2023), VisProg (Gupta & Kembhavi, 2023), LLaVA (Liu et al., 2024), InstructBLIP (Dai et al., 2024), and GPT-4o (OpenAI, 2024). Then, we ablate different components of the DSG framework. Finally, we show that DSG is model-agnostic and can benefit both closed-source and open-source models. Baseline models are given the abstract concept itself in the prompt along with the question (e.g., “Imagine that the image represents a maze.”), but without DSG’s resolved schema of the concept. All results are averaged over $5$ runs. Here, we systematically evaluate multiple-choice answers, where accuracy is calculated as exact-match accuracy. In Appendix A, we provide additional comparisons with a graded accuracy metric based on the degree of alignment to human judgments, and results on free-form answers using BERTScores (Zhang et al., 2020) and LLMs (Kamalloo et al., 2023).
@@
Table 1 summarizes the multiple-choice accuracy of DSG with GPT-4o as the base VLM, in comparison to that of prior works across $5$ runs. DSG significantly outperforms VLMs and integrated LLMs with APIs across all questions and all types of questions. In particular, our method demonstrates a boost in accuracy of $6.6$ percent points overall ($\uparrow 9.9\%$ relative improvement) compared to GPT-4o. Across different question types, DSG shows the largest performance improvement of $10$ percent points ($\uparrow 16.6\%$ relative improvement) for counting-based questions, such as “How many entities are helping?” We hypothesize that the significant improvement in accuracy for these questions is due to more accurate labels for objects with DSG, as the schema grounding step will concretely associate concepts such as “the entity who is helping” with particular parties in the image (e.g., pandas). These questions usually require a more sophisticated reasoning process to answer correctly. In comparison, for binary-choice questions, there may exist more superficial correlations that off-the-shelf VLMs can rely on. Importantly, we also report a random human annotator’s accuracy of adhering to the modal human response to highlight the difficulty of problems in VAB.
In Table 2, we more closely analyze the performance improvement of DSG on GPT-4o across concept categories. We report the average performance and the standard deviation for $5$ runs. We see that DSG outperforms GPT-4o across the four categories. In particular, our method demonstrates a strong boost in accuracy of $9.8$ percent points ($\uparrow 15\%$ relative improvement) for strategic concepts, and $7.6$ percent points ($\uparrow 11.7\%$ relative improvement) for social concepts. In Appendix A, we include qualitative examples of cases where DSG succeeds and fails in correctly grounding schemas. Overall, DSG’s decomposition-based framework improves performance of VLMs across abstract concepts.
Notably, we also analyze the performance of GPT-4v OpenAI (2023) and DSG with GPT-4v as the base model, compared to that of GPT-4o. We summarize results as follows: GPT-4v yields an overall accuracy of $0.653$, GPT-4o yields $0.664$, DSG with GPT-4v yields $0.707$, and DSG with GPT-4o yields $0.730$. We see that DSG significantly improves the performance of both base models. Importantly, DSG with GPT-4v as the base model outperforms GPT-4o by $4.3$ percent points, and showcases the potential for DSG to improve the visual reasoning performance of weaker models.
@@
+ schema $0.681$ $0.629$ $0.672$ $0.709$
+ grounding $0.693$ $0.727$ $0.660$ $0.710$
+ hierarchy $0.701$ $0.713$ $0.671$ $0.723$
DSG (ours) $\mathbf{0.730}$ $\mathbf{0.704}$ $\mathbf{0.690}$ $\mathbf{0.776}$
@@
We present ablation studies of different modules in DSG in Table 3. In particular, we evaluate (1) a variant of DSG that is given only the LLM-generated schema for chain-of-thought reasoning within a single query, without using explicit grounding steps (+ schema), (2) a variant where DSG performs decomposed grounding sequentially instead of hierarchically (+ grounding), and (3) a hierarchical variant (+ hierarchical) where in the final question-answering step, we provide the model with only the grounding result of each schema component, instead of a full history of its grounding process.
With the first variant, we test whether GPT-4o can use the schema appropriately without a structured thinking process. We explore the second variant to test whether decomposing the visual abstraction into a set of key elements is sufficient for improving the understanding of abstract concepts, and evaluate the role of hierarchical dependencies in lower-level concepts for grounding. Finally, we present results on the third variant to test whether the full grounding trace of DSG is helpful during the final question-answering step. The results show that all these design choices help: a schema only does not lend itself to much improved performance, and hierarchical grounding from concrete to abstract symbols, as well as the inclusion of the full context, improves performance in many concepts.
DSG improves other (open-source) VLMs, too.
Tables 4 and 5 summarize the results of DSG with open-source LLaVA as the base VLM. DSG yields consistent improvement overall ($\uparrow 7.0\%$ relative improvement). However, in contrast to GPT-4o, both LLaVA and DSG yield poor results on counting questions, as the base VLM struggles at enumeration regardless of its abstraction understanding.